40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jan 22 02:11:09 2011, kiwamu, Update, Green Locking, some more progress 
    Reply  Sat Jan 22 02:23:26 2011, Koji, Update, Green Locking, some more progress 
Message ID: 4189     Entry time: Sat Jan 22 02:11:09 2011     Reply to this: 4190
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: some more progress 

[Rana, Suresh, Kiwanu]

 We did the following things:

   *  taking the VCO stability data from the error signal instead of the feedback

   *  tried calibrating the signal but confused

   *  increased the modulation depth of the green end PDH.

--

 We found that a cable coming out from the VCO box was quite touchy. This cable was used for taking the feedback signal.

When we touched the cable it made a big noise in the feedback. So we decided to remove the cable and take the signal from the error point (i.e. just after the mixer and the LPF.)

In order to correct that signal to the one in terms of the feedback signal, we put a digital filter which is exactly the same as that of the PLL (pole at 1.5 Hz, zero at 40 Hz, G=1) .

However for some reasons the signal shown in the digital side looked completely mis-calibrated by ~ 100. We have no idea what is going on.

Anyway we are taking the data over tonight because we can correct the signal later. The 2nd round data started from AM1:40

ELOG V3.1.3-